OpenStudy (anonymous):

How much force is needed to stop a 90-kg soccer player if he decelerates at 15 m/s²? A. 1350 N B. 15 N C. 6 N D. 135 N

OpenStudy (anonymous):

force = mass * acceleration

OpenStudy (anonymous):

So how do I do that? @sourwing

OpenStudy (anonymous):

F = m* a m = mass = 90 kg a = acceleration = 15 m/s^2 F = force